If f(x) = 2x-4 And g(x) = x2-4x
1. F(4)
How Do I Do This

Answer by nyc_function(2741)

If f(x) = 2x-4 And g(x) = x^2-4x
Find f(4).

Just replace x with 4 for the function f(x) and simplify.

f(x) = 2x - 4

f(4) = 2(4) - 4

f(4) = 8 - 4

f(4) = 4
Done!!